# Open dynamical systems with a moving hole

### v1

@@ Abstract.
Given an integer $b\geq 3$, let $T_{b}:[0,1)\to[0,1);x\mapsto bx\pmod{1}$ be the expanding map on the unit circle. For any $m\in\mathbb{N}$ and for any $\omega=\omega^{0}\omega^{1}\ldots\in(\left\{0,1,\ldots,b-1\right\}^{m})^{\mathbb{N}}$ let
$K^{\omega}=\left\{x\in[0,1):T_{b}^{n}(x)\notin I_{\omega^{n}}\leavevmode\nobreak\ \forall n\geq 0\right\},$
where $I_{\omega^{n}}$ is the $b$-adic basic interval generated by $\omega^{n}$. Then $K^{\omega}$ is called the survivor set of the open dynamical system $([0,1),T_{b},I_{\omega})$ with respect to the sequence of holes $I_{\omega}=\left\{I_{\omega^{n}}:n\geq 0\right\}$. We show that the Hausdorff and lower box dimensions of $K^{\omega}$ always conincide, and the packing and upper box dimensions of $K^{\omega}$ also coincide. Moreover, we give sharp lower and upper bounds for the dimensions of $K^{\omega}$, which can be calculated explicitly. Finally, for any admissible $\alpha\leq\beta$ we show that there exist infinitely many $\omega$ (in fact of positive dimension) such that $\dim_{H}K^{\omega}=\alpha$ and $\dim_{P}K^{\omega}=\beta$.
@@
Open dynamical system; Hausdorff dimension; packing dimension; badly approximable numbers; finiteness property; joint spectral radius
@@ 1. Introduction
The mathematical study of dynamical systems with holes, called open dynamical systems, was first proposed by Pianigiani and Yorke [43]. Let $(X,T,\mu)$ be an ergodic dynamical system on a compact metric space $X$, and let $H\subset X$ be the hole with $\mu(H)>0$. For $n\in\mathbb{N}$ let $P_{s}(n)$ be the survivor probability defined by $P_{s}(n):=\mu(\left\{x\in X:T^{j}x\notin H\leavevmode\nobreak\ \forall 0\leq j\leq n\right\})$. Open dynamical systems always study the decay rate of $P_{s}(n)$, called the escape rate, given by $\rho(H):=\lim_{n\to\infty}-\frac{1}{n}\log P_{s}(n)$. For more information on the escape rate and the related first hitting time we refer to the papers [9, 17, 18, 19] and the references therein. Note by Birkhoff ergodic theorem (cf. [52]) that the survivor set
$K(H):=\left\{x\in X:T^{n}x\notin H\leavevmode\nobreak\ \forall n\in\mathbb{N}_{0}\right\}$
is a $\mu$-null set. Then the fractal dimension, such as Hausdorff dimension, is applied to estimate the size of $K(H)$. It turns out that the dimension of $K(H)$ depends not only on the size but also the position of $H$ (cf. [11]). For more results on the dimension of $K(H)$ we refer to the papers [4, 27, 32, 50]. Note that the study of $K(H)$ has intimate connection with unique $\beta$-expansions (cf. [2, 3, 16, 26, 33, 46]) and badly approximable numbers in Diophantine approximation (cf. [15, 22, 36, 41, 50]).
Until now, in most work of open dynamical systems the hole $H$ was assumed to be fixed. In this paper we will study the survivor set of an open dynamical system with a moving hole. In particular, the hole moves when the time $n$ changes. Our study involves a new type of fractals including the Moran type fractals and graph-directed fractals as special cases. Our results can be applied to the study of limitinf sets in dynamical Diophantine approximations and to the finiteness property of joint spectral radius of matrices.
@@ 1.1. Open dynamical systems with a moving hole
Let :=$\mathbb{N}:=\{1,2,3,\ldots\}$ be the set of all positive integers, and let $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $k\in\mathbb{N}$ let $\mathbb{N}_{\geq k}:=\mathbb{N}\cap[k,+\infty)$.
Given $b\in\mathbb{N}_{\geq 3}$, let $D_{b}:=\left\{0,1,\ldots,b-1\right\}$. For $n\in\mathbb{N}$ let $D_{b}^{n}$ be the set of all words of length $n$ over the alphabet $D_{b}$, i.e., $D_{b}^{n}=\{d_{1}d_{2}\ldots d_{n}:d_{i}\in D_{b}\leavevmode\nobreak\ \forall 1\leq i\leq n\}.$ Let $D_{b}^{*}:=\bigcup_{n=1}^{\infty}D_{b}^{n}$ be the set of all finite words over the alphabet $D_{b}$. Moreover, let $D_{b}^{\mathbb{N}}:=\{(d_{i})_{i=1}^{\infty}:d_{i}\in D_{b}\leavevmode\nobreak\ \forall i\in\mathbb{N}\}$ denote the set of all infinite sequences over the alphabet $D_{b}$, and let $\sigma$ be the left shift map on $D_{b}^{\mathbb{N}}$. Then for $(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}$ we have $\sigma((d_{i})_{i=1}^{\infty})=(d_{i+1})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}$. For a finite word $\mathbf{d}=d_{1}d_{2}\ldots d_{n}\in D_{b}^{*}$, we define the cylinder set by $\mathcal{C}_{\mathbf{d}}:=\{(\omega_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:\omega_{i}=d_{i}\leavevmode\nobreak\ \forall 1\leq i\leq n\}$.
Given $m\in\mathbb{N}$, write $(D_{b}^{m})^{\mathbb{N}}$ for the set of all infinite sequences $\omega=\omega^{0}\omega^{1}\ldots$ with each $\omega^{k}\in D_{b}^{m}$. For $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ we first define the survivor set of the symbolic dynamical system $(D_{b}^{\mathbb{N}},\sigma)$ with the moving holes $\{\mathcal{C}_{\omega^{k}}:k\in\mathbb{N}_{0}\}$ by
$\Sigma^{\omega}:=\big{\{}(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:\sigma^{k}\big{(}(d_{i})_{i=1}^{\infty}\big{)}\not\in\mathcal{C}_{\omega^{k}}\leavevmode\nobreak\ \forall k\in\mathbb{N}_{0}\big{\}}.$
@@
(1.1) $\Sigma^{\omega}=\big{\{}(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:d_{k+1}d_{k+2}\ldots d_{k+m}\neq\omega^{k}\leavevmode\nobreak\ \forall k\in\mathbb{N}_{0}\big{\}}.$
@@
It is closely related to the survivor set of the open dynamical system $([0,1),T_{b})$ with the moving holes $\big{\{}\mathrm{int}\big{(}\pi_{b}(\mathcal{C}_{\omega^{k}})\big{)}:k\in\mathbb{N}_{0}\}$, where $T_{b}:[0,1)\to[0,1);\leavevmode\nobreak\ x\mapsto bx\pmod{1}$. Observe that up to a countable set the map $\pi_{b}$ is conjugate from $(D_{b}^{\mathbb{N}},\sigma)$ to $([0,1),T_{b})$ (cf. [38]). So, the survivor set
$\widetilde{K}^{\omega}:=\big{\{}x\in[0,1):T_{b}^{k}(x)\not\in\mathrm{int}\big{(}\pi_{b}(\mathcal{C}_{\omega^{k}})\big{)}\leavevmode\nobreak\ \forall k\in\mathbb{N}_{0}\big{\}}$
@@
Since $b\in\mathbb{N}_{\geq 3}$, the survivor set $K^{\omega}$ is a Cantor set, i.e., $K^{\omega}$ is a compact set with neither interior nor isolated points. Our work involves determining the various fractal dimensions of the survivor set $K^{\omega}$.
For a bound set $F\subset\mathbb{R}^{n}$, let $\dim_{H}F,\leavevmode\nobreak\ \dim_{P}F$ denote its Hausdorff and packing dimensions; let $\underline{\dim}_{B}F,\leavevmode\nobreak\ \overline{\dim}_{B}F$ denote its lower and upper box dimensions (cf. [20]); and let $\dim_{L}F,\leavevmode\nobreak\ \dim_{A}F$ denote its lower dimension and Assouad dimension (cf. [25]). For a compact set $F\subset\mathbb{R}^{n}$, these fractal dimensions have the following relation (cf. [25, Lemma 2.4.3, Theorem 3.4.3])
@@
In general, the lower box dimension and the packing dimension are incomparable. If $F$ is $s$-Ahlfors regular, then all fractal dimensions in (1.4) are equal to $s$. Recall that a compact set $F\subset\mathbb{R}^{n}$ is called $s$-Ahlfors regular if there exists a constant $c\geq 1$ such that for all $10<r<$ 1 and all $x\in F$,
@@
When $\omega^{k}=\mathbf{d}$ for all $k\in\mathbb{N}_{0}$, we have $(\Sigma^{\omega},\sigma)$ is the shift of finite type with the forbidden word $\mathbf{d}$, denoted by $X_{\mathbf{d}}$ (cf. [38]). It can be described by a directed graph $G_{\mathbf{d}}=(V,E)$, where $V=D_{b}^{m-1}$ and $(\mathbf{u},\mathbf{v})\in E$ for $\mathbf{u}=u_{1}\ldots u_{m-1},\mathbf{v}=v_{1}\ldots v_{m-1}\in V$ if $u_{2}\ldots u_{m-1}=v_{1}\ldots v_{m-2}$ and $u_{1}\ldots u_{m-1}v_{m-1}\neq\mathbf{d}$. Let $A_{\mathbf{d}}$ be the corresponding adjacency matrix of $G_{\mathbf{d}}$, whose size is $b^{m-1}\times b^{m-1}$. In this case, all fractal dimensions in (1.4) of $K^{\omega}=\pi_{b}(X_{\mathbf{d}})$ are equal to $\log\rho(A_{\mathbf{d}})/\log b$, where $\rho(A)$ is the spectral radius of a matrix $A$ (cf. [40]). To the best of our knowledge, $K^{\omega}$ is a new type of fractal set for a general $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$.
@@
For any $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ we have
@@
By a recent work of Fan and Wu [21] it follows that for any ergodic measure $\nu$ on $((D_{b}^{m})^{\mathbb{N}},\tilde{\sigma})$ we have
$\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\lim_{n\to\infty}\frac{1}{n\log b}\log\|A_{\omega^{0}}A_{\omega^{1}}\cdots A_{\omega^{n-1}}\|\quad\textrm{for $\nu$-a.e. }\omega\in(D_{b}^{m})^{\mathbb{N}},$
where $\tilde{\sigma}(\omega)=\omega^{1}\omega^{2}\ldots$ for $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$.
To describe the range of $\dim_{H}K^{\omega}$ and $\dim_{P}K^{\omega}$ we introduce two types of sequences in $(D_{b}^{m})^{\mathbb{N}}$, namely the progressively overlapping sequences and the totally distinct sequences. Let $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ with $\omega^{i}=\omega^{i}_{1}\ldots\omega^{i}_{m}\in D_{b}^{m}$ for all $i\in\mathbb{N}_{0}$. For $k\in\mathbb{N}$, we say that the sequence $\omega$ is progressively overlapping at position-$k$ if
@@
The sequence $\omega$ is called progressively overlapping if it is progressively overlapping at position-$k$ for all $k\in\mathbb{N}$, that is, $\omega^{i}_{1}\ldots\omega^{i}_{m-1}=\omega^{i-1}_{2}\ldots\omega^{i-1}_{m}$ for all $i\in\mathbb{N}$. Each sequence $d_{1}d_{2}\ldots\in D_{b}^{\mathbb{N}}$ induces a progressively overlapping sequence $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ by letting $\omega^{i}=d_{i+1}\ldots d_{i+m}$ for all $i\in\mathbb{N}_{0}$.
@@
Let $b\in\mathbb{N}_{\geq 3}$ and $m=2$. For any $d_{1}d_{2}d_{3}\ldots\in D_{b}^{\mathbb{N}}$ let $\omega^{i}=d_{i+1}d_{i+2}$ for all $i\in\mathbb{N}_{0}$. Then $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ is a progressively overlapping sequence. Similarly, let $\tau^{i}=c_{i+1}d_{i+2}$ with $c_{i+1}\neq d_{i+1}$ for all $i\in\mathbb{N}_{0}$. Then $\tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ is a totally distinct sequence. So, there exist uncountably many progressively overlapping sequences and uncountably many totally distinct sequences.
When $\omega\in(D_{b}^{m})^{\mathbb{N}}$ is a progressively overlapping or a totally distinct sequence, the dimensions in (1.4) are all equal for $K^{\omega}$.
@@
    If $\omega\in(D_{b}^{m})^{\mathbb{N}}$ is a progressively overlapping sequence, then the set $K^{\omega}$ is $\frac{\log\lambda}{\log b}$-Ahlfors regular, where $\lambda\in(b-1,b)$ is a Pisot number satisfying the equation $x^{m}-(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)=0$.
@@
    If $\omega\in(D_{b}^{m})^{\mathbb{N}}$ is a totally distinct sequence, then the set $K^{\omega}$ is $\frac{\log\eta}{\log b}$-Ahlfors regular, where $\eta\in(b-1,b)$ is a Pisot number satisfying the equation $x^{m}-bx^{m-1}+1=0$.
@@
If $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ is eventually progressively overlapping, i.e., there exists $k_{0}\in\mathbb{N}$ such that $\omega$ is progressively overlapping at position-$k$ for all $k\geq k_{0}$, then the set $K^{\omega}$ is also $\frac{\log\lambda}{\log b}$-Ahlfors regular. Similarly, if $\omega$ is eventually totally distinct, then the set $K^{\omega}$ is $\frac{\log\eta}{\log b}$-Ahlfors regular.
Finally, we show that the dimensions obtained in Theorem 1.4 (i) and (ii) are the upper and lower bounds for fractal dimensions of $K^{\omega}$, respectively; and the Hausdorff and packing dimensions have the intermediate value property. Furthermore, the strict inequalities
@@
hold for infinitely many sequences $\omega\in(D_{b}^{m})^{\mathbb{N}}$. In order to consider the Hausdorff dimension in the symbolic space $(D_{b}^{m})^{\mathbb{N}}$, we equip $(D_{b}^{m})^{\mathbb{N}}$ with the following metric
(1.5) $\varrho(\omega,\tau)=b^{-m\cdot\inf\{k\in\mathbb{N}_{0}:\omega^{k}\neq\tau^{k}\}}\quad\text{for}\quad\omega=\omega^{0}\omega^{1}\ldots,\leavevmode\nobreak\ \tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}.$
@@
    For any $\omega\in(D_{b}^{m})^{\mathbb{N}}$ we have
@@
    $\dim_{H}\Big{\{}\omega\in(D_{b}^{m})^{\mathbb{N}}:\dim_{H}K^{\omega}=\alpha,\leavevmode\nobreak\ \dim_{P}K^{\omega}=\beta\Big{\}}\geq\frac{1}{m}.$
@@
    $\dim_{H}\Big{\{}\omega\in(D_{b}^{m})^{\mathbb{N}}:\dim_{L}K^{\omega}<\dim_{H}K^{\omega}<\dim_{P}K^{\omega}<\dim_{A}K^{\omega}\Big{\}}\geq\frac{1}{m}.$
@@
    $\sum_{k=1}^{N}\left(\sum_{\mathbf{d}\in D_{b}^{m}}A_{\mathbf{d}}\right)^{k}>\bm{0}.$
    So by using Theorem 1.1 of [23] we obtain the following multifractal result of $K^{\omega}$. For any $\alpha\in[\log_{b}\eta,\log_{b}\lambda]$ set $L_{\alpha}:=\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}}:\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\alpha\right\}.$ Then [23, Theorem 1.1] implies that

### v2

@@ Abstract.
Given an integer $b\geq 3$, let $T_{b}:[0,1)\to[0,1);x\mapsto bx\pmod{1}$ be the expanding map on the unit circle. For any $m\in\mathbb{N}$ and for any $\omega=\omega^{0}\omega^{1}\ldots\in(\left\{0,1,\ldots,b-1\right\}^{m})^{\mathbb{N}_{0}}$ let
$K^{\omega}=\left\{x\in[0,1):T_{b}^{n}(x)\notin I_{\omega^{n}}\penalty 10000\ \forall n\geq 0\right\},$
where $I_{\omega^{n}}$ is the $b$-adic basic interval generated by $\omega^{n}$. Then $K^{\omega}$ is called the survivor set of the open dynamical system $([0,1),T_{b},I_{\omega})$ with respect to the sequence of holes $I_{\omega}=\left\{I_{\omega^{n}}:n\geq 0\right\}$. We show that the Hausdorff and lower box dimensions of $K^{\omega}$ always coincide, and the packing and upper box dimensions of $K^{\omega}$ also coincide. Moreover, we give sharp lower and upper bounds for the dimensions of $K^{\omega}$, which can be calculated explicitly. Finally, for any admissible $\alpha\leq\beta$ we show that there exist infinitely many $\omega$ (in fact of positive dimension) such that $\dim_{H}K^{\omega}=\alpha$ and $\dim_{P}K^{\omega}=\beta$.
@@
Open dynamical system; Hausdorff dimension; packing dimension; badly approximable numbers; non-recurrence points; joint spectral radius
@@ 1. Introduction
The mathematical study of dynamical systems with holes, called open dynamical systems, was first proposed by Pianigiani and Yorke [43]. Let $(X,T,\mu)$ be an ergodic dynamical system on a compact metric space $X$, and let $H\subset X$ be the hole with $\mu(H)>0$. For $n\in\mathbb{N}$ let $P_{s}(n)$ be the survivor probability defined by $P_{s}(n):=\mu\left(\left\{x\in X:T^{j}(x)\notin H\penalty 10000\ \forall 0\leq j\leq n\right\}\right)$. In open dynamical systems people always study the decay rate of $P_{s}(n)$, called the escape rate, given by $\rho(H):=\lim_{n\to\infty}-\frac{1}{n}\log P_{s}(n)$. For more information on the escape rate and the related first hitting time we refer to the papers [9, 17, 18, 19] and the references therein. Note by Birkhoff ergodic theorem (cf. [52]) that the survivor set
$K(H):=\left\{x\in X:T^{n}(x)\notin H\penalty 10000\ \forall n\in\mathbb{N}_{0}\right\}$
is a $\mu$-null set. Then the fractal dimension, such as Hausdorff dimension, is applied to estimate the size of $K(H)$. It turns out that the dimension of $K(H)$ depends not only on the size but also on the position of $H$ (cf. [11]). For more results on the dimension of $K(H)$ we refer to the papers [4, 27, 32, 50]. Note that the study of $K(H)$ has intimate connection with unique $\beta$-expansions (cf. [2, 3, 16, 26, 33, 46]) and badly approximable numbers in Diophantine approximation (cf. [15, 22, 36, 41]).
Until now, in most work of open dynamical systems the hole $H$ was assumed to be fixed. In this paper we will study the survivor set of an open dynamical system with a moving hole. In particular, the hole moves when the time $n$ changes. Our study involves a new type of fractals including the Moran type fractals and graph-directed fractals as special cases. Our results can be applied to the study of liminf sets in dynamical Diophantine approximations and to the finiteness property for the joint spectral radius of matrices.
@@ 1.1. Open dynamical systems with a moving hole
Let $\mathbb{N}:=\{1,2,3,\ldots\}$ be the set of all positive integers, and let $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $k\in\mathbb{N}$ let $\mathbb{N}_{\geq k}:=\mathbb{N}\cap[k,+\infty)$.
Given $b\in\mathbb{N}_{\geq 3}$, let $D_{b}:=\left\{0,1,\ldots,b-1\right\}$. For $n\in\mathbb{N}$ let $D_{b}^{n}$ be the set of all words of length $n$ over the alphabet $D_{b}$, i.e., $D_{b}^{n}=\{d_{1}d_{2}\ldots d_{n}:d_{i}\in D_{b}\penalty 10000\ \forall 1\leq i\leq n\}.$ Let $D_{b}^{*}:=\bigcup_{n=1}^{\infty}D_{b}^{n}$ be the set of all finite words over the alphabet $D_{b}$. Moreover, let $D_{b}^{\mathbb{N}}:=\{(d_{i})_{i=1}^{\infty}:d_{i}\in D_{b}\penalty 10000\ \forall i\in\mathbb{N}\}$ denote the set of all infinite sequences over the alphabet $D_{b}$, and let $\sigma$ be the left shift map on $D_{b}^{\mathbb{N}}$. Then for $(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}$ we have $\sigma((d_{i})_{i=1}^{\infty})=(d_{i+1})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}$. For a finite word $\mathbf{d}=d_{1}d_{2}\ldots d_{n}\in D_{b}^{*}$, we define the cylinder set by $\mathcal{C}_{\mathbf{d}}:=\{(\omega_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:\omega_{i}=d_{i}\penalty 10000\ \forall 1\leq i\leq n\}$.
Given $m\in\mathbb{N}$, write $(D_{b}^{m})^{\mathbb{N}_{0}}$ for the set of all infinite sequences $\omega=\omega^{0}\omega^{1}\ldots$ with each $\omega^{k}\in D_{b}^{m}$. For $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ we first define the survivor set of the symbolic dynamical system $(D_{b}^{\mathbb{N}},\sigma)$ with the moving holes $\{\mathcal{C}_{\omega^{k}}:k\in\mathbb{N}_{0}\}$ by
$\Sigma^{\omega}:=\big\{(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:\sigma^{k}\big((d_{i})_{i=1}^{\infty}\big)\not\in\mathcal{C}_{\omega^{k}}\penalty 10000\ \forall k\in\mathbb{N}_{0}\big\}.$
@@
(1.1) $\Sigma^{\omega}=\big\{(d_{i})_{i=1}^{\infty}\in D_{b}^{\mathbb{N}}:d_{k+1}d_{k+2}\ldots d_{k+m}\neq\omega^{k}\penalty 10000\ \forall k\in\mathbb{N}_{0}\big\}.$
@@
It is closely related to the survivor set of the open dynamical system $([0,1),T_{b})$ with the moving holes $\big\{\mathrm{int}\big(\pi_{b}(\mathcal{C}_{\omega^{k}})\big):k\in\mathbb{N}_{0}\}$, where $T_{b}:[0,1)\to[0,1);\penalty 10000\ x\mapsto bx\pmod{1}$. Observe that up to a countable set the map $\pi_{b}$ is conjugate from $(D_{b}^{\mathbb{N}},\sigma)$ to $([0,1),T_{b})$ (cf. [38]). So, the survivor set
$\widetilde{K}^{\omega}:=\big\{x\in[0,1):T_{b}^{k}(x)\not\in\mathrm{int}\big(\pi_{b}(\mathcal{C}_{\omega^{k}})\big)\penalty 10000\ \forall k\in\mathbb{N}_{0}\big\}$
@@
Since $b\in\mathbb{N}_{\geq 3}$, the survivor set $K^{\omega}$ is a Cantor set, i.e., $K^{\omega}$ is a compact set with neither interior nor isolated points. Our work involves determining various fractal dimensions of the survivor set $K^{\omega}$.
For a bounded set $F\subset\mathbb{R}^{n}$, let $\dim_{H}F,\penalty 10000\ \dim_{P}F$ denote its Hausdorff and packing dimensions; let $\underline{\dim}_{B}F,\penalty 10000\ \overline{\dim}_{B}F$ denote its lower and upper box dimensions (cf. [20]); and let $\dim_{L}F,\penalty 10000\ \dim_{A}F$ denote its lower dimension and Assouad dimension (cf. [25]). For a compact set $F\subset\mathbb{R}^{n}$, these fractal dimensions have the following relation (cf. [25, Lemma 2.4.3, Theorem 3.4.3])
@@
In general, the lower box dimension and the packing dimension are incomparable. If $F$ is $s$-Ahlfors regular, then all fractal dimensions in (1.4) are equal to $s$. Recall that a compact set $F\subset\mathbb{R}^{n}$ is called $s$-Ahlfors regular if there exists a constant $c\geq 1$ such that for all $0<r<1$ and all $x\in F$,
@@
When $\omega^{k}=\mathbf{d}$ for all $k\in\mathbb{N}_{0}$, we have $(\Sigma^{\omega},\sigma)$ is the shift of finite type with the forbidden word $\mathbf{d}$, denoted by $X_{\mathbf{d}}$ (cf. [38]). It can be described by a directed graph $G_{\mathbf{d}}=(V,E)$, where $V=D_{b}^{m-1}$, and for $\mathbf{u}=u_{1}\ldots u_{m-1},\mathbf{v}=v_{1}\ldots v_{m-1}\in V$ we say $(\mathbf{u},\mathbf{v})\in E$ if $u_{2}\ldots u_{m-1}=v_{1}\ldots v_{m-2}$ and $u_{1}\ldots u_{m-1}v_{m-1}\neq\mathbf{d}$. Let $A_{\mathbf{d}}$ be the corresponding adjacency matrix of $G_{\mathbf{d}}$, whose size is $b^{m-1}\times b^{m-1}$. In this case, all fractal dimensions in (1.4) of $K^{\omega}=\pi_{b}(X_{\mathbf{d}})$ are equal to $\log\rho(A_{\mathbf{d}})/\log b$, where $\rho(A)$ is the spectral radius of a matrix $A$ (cf. [40]). To the best of our knowledge, $K^{\omega}$ is a new type of fractal set for a general $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$.
@@
For any $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ we have
@@
[…]
$\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\lim_{n\to\infty}\frac{1}{n\log b}\log\|A_{\omega^{0}}A_{\omega^{1}}\cdots A_{\omega^{n-1}}\|\quad\textrm{for $\nu$-a.e. }\omega\in(D_{b}^{m})^{\mathbb{N}_{0}},$
where $\tilde{\sigma}(\omega)=\omega^{1}\omega^{2}\ldots$ for $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$.
To describe the range of $\dim_{H}K^{\omega}$ and $\dim_{P}K^{\omega}$ we introduce two types of sequences in $(D_{b}^{m})^{\mathbb{N}_{0}}$, namely the progressively overlapping sequences and the totally distinct sequences. Let $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ with $\omega^{i}=\omega^{i}_{1}\ldots\omega^{i}_{m}\in D_{b}^{m}$ for all $i\in\mathbb{N}_{0}$. For $k\in\mathbb{N}$, we say that the sequence $\omega$ is progressively overlapping at position-$k$ if
@@
The sequence $\omega$ is called progressively overlapping if it is progressively overlapping at position-$k$ for all $k\in\mathbb{N}$, that is, $\omega^{i}_{1}\ldots\omega^{i}_{m-1}=\omega^{i-1}_{2}\ldots\omega^{i-1}_{m}$ for all $i\in\mathbb{N}$. Each sequence $d_{1}d_{2}\ldots\in D_{b}^{\mathbb{N}}$ induces a progressively overlapping sequence $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ by letting $\omega^{i}=d_{i+1}\ldots d_{i+m}$ for all $i\in\mathbb{N}_{0}$.
@@
Let $b\in\mathbb{N}_{\geq 3}$ and $m=2$. For any $d_{1}d_{2}d_{3}\ldots\in D_{b}^{\mathbb{N}}$ let $\omega^{i}=d_{i+1}d_{i+2}$ for all $i\in\mathbb{N}_{0}$. Then $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is a progressively overlapping sequence. Similarly, let $\tau^{i}=c_{i+1}d_{i+2}$ with $c_{i+1}\neq d_{i+1}$ for all $i\in\mathbb{N}_{0}$. Then $\tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is a totally distinct sequence. So, there exist uncountably many progressively overlapping sequences and uncountably many totally distinct sequences.
When $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is a progressively overlapping or a totally distinct sequence, the dimensions in (1.4) are all equal for $K^{\omega}$.
@@
    If $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is a progressively overlapping sequence, then the set $K^{\omega}$ is $\frac{\log\lambda}{\log b}$-Ahlfors regular, where $\lambda\in(b-1,b)$ is Pisot number satisfying the equation $x^{m}-(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)=0$.
@@
    If $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is a totally distinct sequence, then the set $K^{\omega}$ is $\frac{\log\eta}{\log b}$-Ahlfors regular, where $\eta\in(b-1,b)$ is the Pisot number satisfying the equation $x^{m}-bx^{m-1}+1=0$.
@@
If $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ is eventually progressively overlapping, i.e., there exists $k_{0}\in\mathbb{N}$ such that $\omega$ is progressively overlapping at position-$k$ for all $k\geq k_{0}$, then the set $K^{\omega}$ is also $\frac{\log\lambda}{\log b}$-Ahlfors regular. Similarly, if $\omega$ is eventually totally distinct, then the set $K^{\omega}$ is $\frac{\log\eta}{\log b}$-Ahlfors regular.
Next, we show that the dimensions obtained in Theorem 1.4 (i) and (ii) are the upper and lower bounds for fractal dimensions of $K^{\omega}$, respectively; and the Hausdorff and packing dimensions have the intermediate value property. Furthermore, the strict inequalities
@@
hold for infinitely many sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$. In order to consider the Hausdorff dimension in the symbolic space $(D_{b}^{m})^{\mathbb{N}_{0}}$, we equip $(D_{b}^{m})^{\mathbb{N}_{0}}$ with the following metric
(1.5) $\varrho(\omega,\tau)=b^{-m\cdot\inf\{k\in\mathbb{N}_{0}:\omega^{k}\neq\tau^{k}\}}\quad\text{for}\quad\omega=\omega^{0}\omega^{1}\ldots,\penalty 10000\ \tau=\tau^{0}\tau^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}.$
@@
    For any $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ we have
@@
    $\dim_{H}\Big\{\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}:\dim_{H}K^{\omega}=\alpha,\penalty 10000\ \dim_{P}K^{\omega}=\beta\Big\}\geq\frac{1}{m}.$
@@
    $\dim_{H}\Big\{\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}:\dim_{L}K^{\omega}<\dim_{H}K^{\omega}<\dim_{P}K^{\omega}<\dim_{A}K^{\omega}\Big\}\geq\frac{1}{m}.$
@@
    $\sum_{k=1}^{N}\left(\sum_{\mathbf{d}\in D_{b}^{m}}A_{\mathbf{d}}\right)^{k}>\bm{0},$
    which means each element in the matrix $\sum_{k=1}^{N}\left(\sum_{\mathbf{d}\in D_{b}^{m}}A_{\mathbf{d}}\right)^{k}$ is positive. So by using Theorem 1.1 of [23] we obtain the following multifractal result of $K^{\omega}$. For any $\alpha\in[\log_{b}\eta,\log_{b}\lambda]$ set $L_{\alpha}:=\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}:\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\alpha\right\}.$ Then [23, Theorem 1.1] implies that